# Constraints on inflationary decoherence from attractor model

### v1

@@ III Models and numerical results
We choose the modified polynomial attractor model in [45]
@@
where $f(\psi)=1/\psi^{3}$, the parameters of the model are given by Table. 1.
From Eq. (8) and (15), we can see that for linear interaction,
@@
$-2v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{R}}_{\bm{k}}(\eta)v^{\mathrm{I}}_{\bm{k}}(\eta),$ (34)
We take $m=3$ so that $\mu$ in Eq. (15) vanishes. In order to compare with [32], we take $k_{\gamma}=5\times 10^{-5}\text{Mpc}^{-1}$ and $H_{*}\ell_{E}=10^{-3}$. The top hat function in Eq. (III) can be written as
@@
Then we can calculate the modified power spectrum numerically.
Models $V_{0}$ $\delta$ $c_{1}$ $c_{4}$ $\psi_{*}$ $N$ $n_{s}$ $r$ $k_{\text{peak}}$
A1 8.33×$\times 10^{-10}$ 1.1007×$\times 10^{-3}$1.07 3.2×$\times 10^{-4}$ 3.75 63.2 0.9658 0.025 $3.0\times 10^{8}$
Table 1: Parameters of the model A1.
Figure 1: The black dashed line is the modified power spectrum, the red dotted line is the slow-roll result of $\Delta P_{k}$, the blue solid line is the numerical result of $\Delta P_{k}$ calculated from model A1.
We calculate the modified term $\Delta P_{k}$ of the attractor model A1. The results of the modified power spectrum and $\Delta P_{k}$ are shown in Fig. 1. From Fig. 1, we see that at the pivot scale $k_{*}$, both the slow-roll model and ultra-slow-roll model have the same value of $\Delta P_{k}$. However, as the Hubble parameter gradually decreases, $\Delta P_{k}$ also decreases. We also find that at the minimum of the power spectrum, $\Delta P_{k}$ has a peak in the order of $10^{-6}$, but the correction near the peak is negligible ($\sim 10^{-8}$). This means that the correction due to decoherence almost does not change the abundance of primordial black holes and the scalar induced gravitational waves.
Figure 2: $n_{s}$ and $r$ with respect to $k_{\gamma}$. The red solid lines is the numerical results from model A1, the black dashed lines is the analytical results for slow-roll.
Figure 3: The spectral index and tensor-to-scalar ratio vary with $k_{\gamma}$. The red solid line is the numerical result, the black dashed line is the analytical result, the black dash dot line is the restriction of $n_{s}$ and $r$ at $1\sigma$ and $2\sigma$ from Planck18. As $k_{\gamma}$ increases, the spectral index and tensor-to-scalar ratio decrease.
Figure 4: Decoherence parameter $\delta_{\bm{k}}$ as a function of e-folds $N$ at $k=k_{*}$. The black dashed line is the result of slow-roll model, the red dotted line is the result of model A1. The vertical solid line indicates the e-folds value at the starting point of $\delta_{k}$ evolution.
From Fig. 1, we find that the correction at pivot scale is more significant, which means that we can restrict $k_{\gamma}$ using the spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$ at pivot scale. The analytical results for $n_{s}$ and $r$ in the case of linear interaction are [32]
@@
$r=\frac{r|_{\mathrm{standard}}}{1+\alpha},\quad\alpha=\frac{\pi}{6}\frac{k_{\gamma}^{2}}{k_{*}^{2}}.$ (37)
We perform numerical computations to determine the relationships between $k_{\gamma}$, $n_{s}$ and $r$. The results are shown in Fig. 2. We can see that when $k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$, the results begin to change significantly, when $k_{\gamma}>0.2\text{Mpc}^{-1}$, the changes in the results are minimal. The tensor-to-scalar ratios are well-aligned, whereas the spectral tilts deviate at large $k_{\gamma}$. We plot the results in the $n_{s}-r$ plane to restrict the upper bound of $k_{\gamma}$, the results are shown in Fig. 3. Starting from the initial point, as $k_{\gamma}$ grows, both $n_{s}$ and $r$ diminish. When $k_{\gamma}=0.026\,\text{Mpc}^{-1}$ and $k_{\gamma}=0.042\,\text{Mpc}^{-1}$, the results are within the $1\sigma$ and $2\sigma$ confidence intervals of the observation, respectively. This is the upper bound of $k_{\gamma}$ given by model A1.
We then calculate the decoherence parameter $\delta_{\bm{k}}$ at pivot scale and compare with the result in [32]. The results are shown in Fig. 4. Since model A1 is in the slow-roll stage at the pivot scale, the analytical and numerical results are almost the same. After $N-N_{*}=4$, $\delta_{\bm{k}}$ increase exponentially. We require that for the scales $k_{*}$ and $k_{\mathrm{peak}}$, decoherence is complete. For model A1, this imposes a constraint on $k_{\gamma}$ such that $k_{\gamma}$ must be much greater than $10^{-28}\,\text{Mpc}^{-1}$. This is a rather relaxed constraint.
@@ IV Conclusion
In this paper, we examine the effects of decoherence on the power spectrum of a specific inflationary model that incorporates an ultra-slow-roll phase. Decoherence arises from the interaction of a quantum system with its environment and is described by the Lindblad equation. Based on the solutions to the Lindblad equation for linear interactions provided by J. Martin et al. [32], we numerically computed the corrections to the power spectrum of a polynomial attractor model due to quantum decoherence and constrained the interaction parameters $k_{\gamma}$ using observations. We found that decoherence significantly alters the power spectrum only on large scales, and it is consistent with previous work at the pivot scales. As the scale decreases, the corrections to the power spectrum due to decoherence also diminish, and at the trough of the power spectrum, the corrections will result in a peak. We also calculated the relationship between the spectral index and the tensor-to-scalar ratio at the CMB with $k_{\gamma}$, and using the observational data from Planck 2018, we determined an upper limit for $k_{\gamma}$ of $0.042$  Mpc-1. We calculate the decoherence parameter $\delta_{k}$, which is used to characterize the degree of completion of decoherence; decoherence is considered to be complete when $\delta_{k}\gg 1$. We computed the evolution of $\delta_{k}$ over time at the pivot scale, and the results are roughly consistent with the analysis of the slow-roll regime. We require that decoherence be completed at the peak scale of the power spectrum by the end of inflation, which will provide a lower limit for the interaction parameter $k_{\gamma}$. Our model suggests a lower limit of $10^{-28}\,\text{Mpc}^{-1}$. Since exact solutions are only provided for linear interactions, the aforementioned numerical calculations must satisfy the conditions of linear interaction.

### v2

@@ III Models and numerical results
We choose the modified polynomial attractor model with the potential [47]
@@
where $f(\psi)=1/\psi^{3}$, and the parameters of the model are given in Table. 1.
From Eqs. (8) and (15), we see that for linear interaction,
@@
$-2v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{R}}_{\bm{k}}(\eta)v^{\mathrm{I}}_{\bm{k}}(\eta).$ (34)
To simplify the calculation, we take $m=3$ so that $\gamma$ defined in Eq. (15) is independent of $\mu$. In order to make comparison with the results in [31], we take $k_{\gamma}=5\times 10^{-5}\text{ Mpc}^{-1}$ and $H_{*}\ell_{E}=10^{-3}$. The top hat function in Eq. (III) is
@@
With these chosen parameters, we calculate the modified power spectrum numerically and the results are shown in Fig. 1.
Model $V_{0}$ $\delta$ $c_{1}$ $c_{4}$ $\psi_{*}$ $N$ $n_{s}$ $r$ $k_{\text{peak}}$
A 5.3×$\times 10^{-10}$ 9.81×$\times 10^{-4}$1.1 2.3×$\times 10^{-4}$ 4 51.4 0.9743 0.017 $8.7\times 10^{10}$
Table 1: Parameters of the model A.
Figure 1: The power spectrum for the polynomial attractor model A. The black solid line is the modified power spectrum, the red dotted line is the correction to the power spectrum $\Delta P_{k}$ within slow-roll approximation, and the black dashed line is the numerical result for the correction to the power spectrum $\Delta P_{k}$.
From Fig. 1, we see that at the pivot scale $k_{*}$, the numerical result of $\Delta P_{k}$ for the ultra-slow-roll model agrees with analytical one within the slow-roll approximation. This is because the pivot scale crosses the horizon during the slow-roll stage. However, as the Hubble parameter gradually decreases, $\Delta P_{k}$ also decreases. At the minimum of the power spectrum, $\Delta P_{k}$ exhibits a peak on the order of $10^{-4}$; yet, the correction near this peak remains negligible (approximately $\sim 10^{-8}$). This indicates that the decoherence-induced correction has an almost negligible effect on the abundance of primordial black holes and the scalar-induced secondary gravitational waves.
Figure 2: The behavior of the observables $n_{s}$ and $r$ with respect to $k_{\gamma}$. The red solid lines are the numerical results and the black dashed lines are the analytical results within slow-roll approximation.
Figure 3: The spectral index and tensor-to-scalar ratio vary with $k_{\gamma}$. The red solid line represents the numerical results, the black dashed line corresponds to the analytical results within slow-roll approximation, and the black solid lines indicate the $1\sigma$ and $2\sigma$ constraints from P-ACT-LB-BK18. As shown by the lines, $n_{s}$ and $r$ decrease as $k_{\gamma}$ increases.
Figure 4: Decoherence parameter $\delta_{\bm{k}}$ at $k=k_{*}$ as a function of $N-N_{*}$. The red dotted line represents the numerical results, and the black dashed line corresponds to the analytical results within slow-roll approximation. The vertical solid line marks the e-fold value at which the evolution of $\delta_{k}$ begins.
Now we restrict $k_{\gamma}$ using observational constraints on the spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$ at the pivot scale. The analytical results for $n_{s}$ and $r$ in the case of linear interaction within slow-roll approximation are [31]
@@
$r=\frac{r|_{\mathrm{standard}}}{1+\alpha},$ (37)
where $\alpha=\pi k_{\gamma}^{2}/(6k_{*}^{2})$. By varying $k_{\gamma}$, we numerically calculate $n_{s}$ and $r$ at the pivotal scale and the results are shown in Fig. 2. In the range $0.2\,\text{Mpc}^{-1}>k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$, both $n_{s}$ and $r$ vary significantly; outside this range, they remain nearly constant. For the tensor-to-scalar ratio $r$, the numerical result agrees with the analytical one within slow-roll approximation. However, for the scalar spectral tilt $n_{s}$, deviations appear when $k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$. In Fig. 3, we plot the results in $n_{s}-r$ plane along with the $1\sigma$ and $2\sigma$ constraints by P-ACT-LB-BK18. From Fig. 3, it is evident that both $n_{s}$ and $r$ decrease as $k_{\gamma}$ increases. The $1\sigma$ and $2\sigma$ constraints from P-ACT-LB-BK18 give upper bounds $k_{\gamma}<0.041\,\text{Mpc}^{-1}$ and $k_{\gamma}<0.061\,\text{Mpc}^{-1}$, respectively.
To constrain $k_{\gamma}$ using decoherence, we present the results for the decoherence parameter $\delta_{\bm{k}}$ at the pivot scale in Fig. 4. Since the pivotal scale crosses the horizon during the slow-roll phase, the numerical results agree well with the analytical approximations. When $N-N_{*}\geq 4$, where $N$ is the number of e-folds before the end of inflation and $N_{*}$ is the number of e-folds remaining at the horizon crossing of the pivot scale, $\delta_{\bm{k}}$ increases exponentially. We require decoherence to be complete for both the scales $k_{*}$ and $k_{\mathrm{peak}}$ by the end of inflation, leading to the lower bound $k_{\gamma}\gg 10^{-17}\,\text{Mpc}^{-1}$. Compared to the results obtained in Ref. [31], our bound is tighter because we demand that perturbations not only at CMB scales but also at the peak of the power spectrum fully decohere by the end of inflation.
@@ IV Conclusion
In this paper, we examine the effects of decoherence on the power spectrum in an inflationary model featuring an ultra-slow-roll phase. Decoherence, arising from interactions between the quantum system and its environment, is described by the Lindblad equation. Previous studies, employing the slow-roll approximation, calculated the interaction’s impact on the power spectrum and placed preliminary constraints on both the interaction parameter $k_{\gamma}$ and decoherence parameter $\delta_{\bm{k}}$.
Using solutions to the Lindblad equation for linear interactions from Ref. [31], we numerically compute the corrections to the power spectrum of a polynomial attractor model and constrain the interaction parameter $k_{\gamma}$ using observational data. We find that decoherence significantly modifies the power spectrum primarily on large scales, consistent with analytical results obtained under the slow-roll approximation. As the scale decreases, the corrections become smaller, but a peak appears at the trough of the power spectrum.
By analyzing the impact of $k_{\gamma}$ on observables such as the scalar spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$, we establish an upper limit of $k_{\gamma}<0.061$  Mpc-1 from the P-ACT-LB-BK18 data. We also calculate the decoherence parameter $\delta_{k}$ at the pivot scale, which is used to characterize the degree of completion of decoherence. Requiring complete decoherence at both the pivot scale $k_{*}$ and the peak scale $k_{\mathrm{peak}}$ by the end of inflation leads to a tighter lower bound, $k_{\gamma}\gg 10^{-17}\,\text{Mpc}^{-1}$.
In conclusion, we constrain the interaction parameter within the range $10^{-17}\text{Mpc}^{-1}<k_{\gamma}<0.061\text{Mpc}^{-1}$.
Acknowledgements.
This work is supported in part by the National Key Research and Development Program of China under Grant No. 2020YFC2201504.